# Assessing (im)balance in signed brain networks

### v1

@@ II.1 Local properties and dependencies of time series
In analogy with the quantities considered for the study of bipartite signed networks, let us define the total number of positive values as
$B^{+}\equiv\sum_{i=1}^{N}\sum_{t=1}^{T}b_{it}^{+}=\sum_{i=1}^{N}k_{i}^{+}=\sum_{t=1}^{T}\kappa_{t}^{+};$ (25)
the expression above suggests the definition of two (sets of) local positive quantities, i.e.
@@
the positive degree of series $i$, capturing the total number of positive entries characterising the $i$-th series and providing information about its auto-correlation, and
@@
the positive degree at time $t$, capturing the total number of positive entries characterising the $t$-th time step and providing information about the correlations between the members of the whole set of time series; analogously,
@@
with obvious meaning of the symbols. The advantage of adopting Iverson’s brackets is that of ensuring that each quantity is, now, computed on a matrix with positive entries, i.e. is positive as well.
@@ II.2 Concordant and discordant temporal motifs
Any two nodes whose associated time series exhibit a significantly large number of concordant motifs will be connected by a +1+1+ 1, while any two nodes whose associated time series exhibit a significantly large number of discordant motifs will be connected by a 1-1- 1.
Lastly, we would like to stress that our projection scheme is applicable only in the absence of null values - a condition that is naturally satisfied by BOLD time series, representing the focus of this study.
@@ II.3 A scheme for the validated projection of multivariate time series
Let us, now, describe each step in detail.
@@ III Computing the similarity of any
The first step of our method prescribes to measure the degree of similarity of series $i$ and $j$. To this aim, let us consider the quantity named signature Gallo et al. (2024a), defined as the scalar product of the two series: upon defining $\mathbf{b}_{i}$ as the $i$-th row of the matrix $\mathbf{B}$, the signature reads
@@ IV Testing the similarity of any
The second step of our method prescribes to evaluate the statistical significance of the similarity of any two series. To this aim, let us, first, notice that
$-T\leq S_{ij}\leq T$ (37)
where $S_{ij}=-T$ if $C_{ij}=0$ (i.e. if each dyadic motif is composed by a 1-1- 1 and a +1+1+ 1) and $S_{ij}=T$ if $D_{ij}=0$ (i.e. if each dyadic motif is composed by either two 1-1- 1s or two +1+1+ 1s) and, then, clarify that we will employ ERGs defined by linear constraints, hence treat links as independent random variables.
@@ IV.1 Homogeneous benchmark
Let us, now, notice that $S_{ij}$ is a sum of i.i.d. Bernoulli random variables [^13] The finite scheme obeyed by $C_{ijt}-D_{ijt}$ is the same finite scheme obeyed by $b_{it}b_{jt}$., the outcomes of each elementary event being, in fact, 1-1- 1 and +1+1+ 1 instead of 0 and +1+1+ 1. Such a consideration allows us to find the probability distribution obeyed by $S_{ij}$ quite straightforwardly, by starting from the one describing $C_{ij}$ alone, i.e.
@@
[…]
@@ IV.2 Heterogeneous benchmark
and so on. In words, if nodes $i$ and $j$ establish $k=0$ concordant motifs, they establish $T$ discordant motifs and the signature reads $s=-T$; if nodes $i$ and $j$ establish $k=1$ concordant motif, they establish $T-1$ discordant motifs and the signature reads $s=2-T$; if nodes $i$ and $j$ establish $k=2$ concordant motif, they establish $T-2$ discordant motifs and the signature reads $s=4-T$. The same line of reasoning can be repeated for all the admissible values of $k$.
@@ IV.3 The statistical significance of similarity:
The first question can be answered upon calculating the two-sided p-value reading
@@ V Validating the projection of a multivariate time series
[…]
@@
[…]
with $t$ representing the usual single-test significance level, hereby set to the value $t=0.05$: since $H_{i}$ is the hypothesis that the distribution of dyadic motifs established by the $i$-th pair of nodes follows a binomial/Poisson-binomial, rejecting it amounts at connecting the corresponding nodes in the projection. Notice that $|H|=N(N-1)/2$.
@@ VI Benchmarks for the analysis of multivariate time series
In order to evaluate the statistical significance of the similarity of any two series $i$ and $j$, a benchmark is needed: a natural choice leads to adopt the class of null models named Exponential Random Graphs (ERGs). Within such a framework, the constrained maximisation of Shannon entropy
@@
leads to assign the generic bipartite network $\mathbf{B}$ the probability
@@
whose value is determined by the vector $\bm{C}(\mathbf{B})$ of topological constraints via the Hamiltonian $H(\bm{\theta},\bm{C}(\mathbf{B}))=\bm{\theta}\cdot\bm{C}(\mathbf{B})=\sum_{i}\theta_{i}C_{i}(\mathbf{B})$. The use of linear constraints allows $P(\mathbf{B})$ to be written in a factorised form, i.e. as the product of pair-specific probability coefficients Gallo et al. (2025).
@@
Let us, now, generalise such a framework to accommodate models for studying multivariate time series: to this aim, we will follow Marcaccioli and Livan (2020); Gallo et al. (2024a) which, in turn, builds upon the analytical approach introduced in Park and Newman (2004) and further developed in Squartini and Garlaschelli (2011).
Let us, now, make a general observation: as each of the $N$ time series is defined by $T$ time steps, in corresponding of which either the value 1-1- 1 or the value +1+1+ 1 can be assumed, the ensemble is constituted by $|\mathbb{B}|=2^{N\cdot T}$ (bi)adjacency matrices, representing all possible, binary, signed, multivariate time series.
@@ VI.1 Signed Random Graph Model
i.e. by the two global constraints $B^{+}(\mathbf{B})$ and $B^{-}(\mathbf{B})$. Although being formally analogous to the BiSRGM Gallo et al. (2024a), let us name it bSRGM.
According to the bSRGM, each entry of a bipartite signed network is a random variable whose behaviour is described by the coefficients
@@
In order to employ the bSRGM for studying real-world multivariate binary time series, the parameters that define it need to be properly tuned: more specifically, one needs to ensure
@@
with the symbol $\mathbf{B}^{*}$ indicating the empirical multivariate time series under analysis. To this aim, one can maximise the likelihood function $\mathcal{L}_{\text{bSRGM}}(\alpha,\gamma)\equiv\ln P_{\text{bSRGM}}(\mathbf{B}^{*}|\alpha,\gamma)$, where $P_{\text{bSRGM}}(\mathbf{B}^{*}|\alpha,\gamma)$ is defined as in eq. 111, with respect to the unknown parameters that define it Garlaschelli and Loffredo (2008). Such a recipe leads us to find
@@ VI.2 Signed Configuration Model
i.e. by the $2(N+T)$ local constraints $\{k_{i}^{+}(\mathbf{B})\}_{i=1}^{N}$, $\{k_{i}^{-}(\mathbf{B})\}_{i=1}^{N}$, $\{\kappa_{t}^{+}(\mathbf{B})\}_{t=1}^{T}$ and $\{\kappa_{t}^{-}(\mathbf{B})\}_{t=1}^{T}$. Although being formally analogous to the BiSCM Gallo et al. (2024a), let us name it bSCM.
According to the bSCM, each entry of a bipartite signed network is a random variable whose behaviour is described by the coefficients
@@
in other words, $b_{it}$ obeys a generalised Bernoulli distribution whose probability coefficients are determined by the (Lagrange multipliers of the) imposed constraints: the series $i$, at time $t$ assumes the value +1+1+ 1 with probability $p_{it}^{+}$ and the value 1-1- 1 with probability $p_{it}^{-}$.
[…]
@@
let us maximise the likelihood function $\mathcal{L}_{\text{bSCM}}(\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})\equiv\ln P_{\text{bSCM}}(\mathbf{B}^{*}|\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})$, where $P_{\text{bSCM}}(\mathbf{B}^{*}|\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})$ is defined as in eq. A, with respect to the unknown parameters that define it Garlaschelli and Loffredo (2008) (see also Appendix A). Such a recipe leads us to find
@@
a system that can be solved only numerically, along the guidelines provided in Vallarano et al. (2021) (see also Appendix B).
@@ VII Dataset description
[…]
Figure 4: Distributions of the signed, bipartite and monopartite, connectance. Left panel: distributions of the positive and negative bipartite connectance, respectively defined as $c^{+}=B^{+}/(N\cdot T)$ (in blue) and $c^{-}=B^{-}/(N\cdot T)$ (in red), across our 100100100 subjects. Middle and right panel: distributions of the positive and negative monopartite connectance, respectively defined as $\rho^{+}=2L^{+}/N(N-1)$ (in blue) and $\rho^{-}=2L^{-}/N(N-1)$ (in red), across the three, different projections of our 100100100 subjects. Kernel density estimation has been employed for better visualisation; bin widths are computed according to the Freedman-Diaconis rule. How can a large number of bipartite, negative links co-exist with the large number of positive links characterising the naïve projection? Bipartite, negative links are evidently arranged into concordant motifs that, in turn, induce a naïve projection populated by (a majority of) positive links. Notice that the naïve and bSCM-induced distributions overlap, in the right panel.
@@ VIII Results
Let us, now, employ the proposed validation procedure to carry out a ‘binary-to-binary’ analysis, i.e. obtain statistically validated binary projections of multivariate binary time series, defined by the ‘rule’ that any two nodes sharing a significantly large number of concordant (discordant) motifs are connected by a positive (negative) edge (see also fig. 2).

### v3

@@ II.1 Local properties and dependencies of time series
Let us start by defining a number of quantities of interest for our study. In analogy with the properties traditionally considered to analyse bipartite signed networks, the total number of positive values reads
$B^{+}\equiv\sum_{i=1}^{N}\sum_{t=1}^{T}b_{it}^{+}=\sum_{i=1}^{N}k_{i}^{+}=\sum_{t=1}^{T}\kappa_{t}^{+},$ (25)
an expression further suggesting the definition of two (sets of) local positive quantities, i.e.
@@
or positive degree of series $i$, capturing the total number of positive entries characterising the $i$-th series and providing information about its auto-correlation, and
@@
or positive degree at time $t$, capturing the total number of positive entries characterising the $t$-th time step and providing information about the correlations between the members of the whole set of time series. Analogously, the total number of negative values reads
@@
with obvious meaning of the symbols. The density of positive and negative values, reading $c^{+}\equiv B^{+}/(N\cdot T)$ and $c^{-}\equiv B^{-}/(N\cdot T)$ and referred to as positive and negative bipartite connectance, remain naturally defined.
@@ II.2 Concordant and discordant temporal motifs
Any two nodes whose associated time series exhibit a significantly large number of concordant motifs will be connected by a +1+1, while any two nodes whose associated time series exhibit a significantly large number of discordant motifs will be connected by a 1-1. We would also like to stress that our projection scheme is applicable only in the absence of null values - a condition that is naturally satisfied by BOLD time series, representing the focus of this study.
@@ II.3 A scheme for the validated projection of multivariate time series
Let us now describe each step in detail.
@@ III Computing the similarity of any
The first step of our method prescribes measuring the degree of similarity of series $i$ and $j$. To this aim, let us consider the quantity named signature gallo2024testing, defined as the scalar product of the two series: upon defining $\mathbf{b}_{i}$ as the $i$-th row of the matrix $\mathbf{B}$, the signature reads
@@ IV Testing the similarity of any
The second step of our method prescribes evaluating the statistical significance of the similarity of any two series. To this aim, let us first notice that
$T\leq S_{ij}\leq T$ (37)
where $S_{ij}=-T$ if $C_{ij}=0$ (i.e. if each dyadic motif is composed by a 1-1 and a +1+1) and $S_{ij}=T$ if $D_{ij}=0$ (i.e. if each dyadic motif is composed by either two 1-1s or two +1+1s) and, then, clarify that we will employ ERGs defined by linear constraints, hence treat links as independent random variables.
@@ IV.1 Homogeneous benchmark
Let us, now, notice that $S_{ij}$ is a sum of i.i.d. Bernoulli random variables [^13] The finite scheme obeyed by $C_{ijt}-D_{ijt}$ is the same finite scheme obeyed by $b_{it}b_{jt}$., the outcomes of each elementary event being, in fact, 1-1 and +1+1 instead of 0 and +1+1. Such a consideration allows us to find the probability distribution obeyed by $S_{ij}$ quite straightforwardly, by starting from the one describing $C_{ij}$ alone, i.e.
@@
that ranges from T-T to $T$. Let us, now, convince ourselves that this transformation is indeed the correct one, by providing some explicit examples: if nodes $i$ and $j$ establish $k=0$ concordant motifs, they establish $T$ discordant motifs and the signature reads s=Ts=-T; if nodes $i$ and $j$ establish $k=T/2$ concordant motifs, they establish the same number of discordant motifs and the signature reads $s=0$; if nodes $i$ and $j$ establish $k=T$ concordant motifs, they establish zero discordant motifs and the signature reads $s=T$. In other words, the aforementioned change of variable transforms a discrete distribution whose support extends from 0 to $T$ into a discrete distribution whose support extends from T-T to $T$: more intuitively, we have moved from considering sequences of 0s and +1+1s to considering sequences of 1-1s and +1+1s, where each +1+1 can be identified with a concordant motif and each 1-1 can be identified with a discordant motif.
@@ IV.2 Heterogeneous benchmark
and so on. In words, if nodes $i$ and $j$ establish $k=0$ concordant motifs, they establish $T$ discordant motifs and the signature reads s=Ts=-T; if nodes $i$ and $j$ establish $k=1$ concordant motif, they establish T1T-1 discordant motifs and the signature reads s=2Ts=2-T; if nodes $i$ and $j$ establish $k=2$ concordant motif, they establish T2T-2 discordant motifs and the signature reads s=4Ts=4-T. The same line of reasoning can be repeated for all the admissible values of $k$.
@@ IV.3 The statistical significance of similarity:
The first question can be answered upon calculating the two-sided p-value, reading
@@ V Validating the projection of a multivariate time series
One approach is that of adopting the ‘standard’ threshold $t=0.05$ and carrying out each test independently from the others but such a procedure leads to increase [^15] A simple argument can, indeed, be provided. The probability that, solely by chance, at least one out of $|H|$ hypotheses is rejected, and the corresponding link validated, amounts to $\text{FWER}=1-(1-t)^{|H|}$: hence, $\text{FWER}\simeq 1$ for (just) $|H|=100$ tests conducted at the significance level of $t=0.05$. the number of incorrectly rejected null hypotheses (i.e. validated links). An alternative approach is that of adopting the Bonferroni correction BonferroniCarloEmilio1936Tsdc; mantegna2011statistically; gualdi2016statistically but such a procedure is known to severely increase [^16] Indeed, the stricter condition $\text{FWER}=0.05$ leads to the threshold $p_{th}=t/|H|$, rapidly vanishing as $|H|$ grows: as a consequence, very sparse (if not empty) projections are obtained. the number of incorrectly retained null hypotheses (i.e. discarded links).
In the present paper we apply the so-called False Discovery Rate (FDR) procedure benjamini1995controlling. FDR allows one to control for the expected number of ‘false discoveries’ (i.e. incorrectly rejected null hypotheses or incorrectly validated links), regardless of the independence of the hypotheses tested. Whenever $|H|$ different hypotheses $H_{1},\>H_{2}\dots$, characterised by $|H|$ different p-values, must be tested at the same time, FDR prescribes to first sort the $|H|$ p-values in increasing order
$\text{p-value}_{1}\leq\dots\leq\text{p-value}_{|H|}$ (57)
@@
$\text{p-value}_{\hat{i}}\leq\frac{\hat{i}t}{|H|}$ (58)
with $t$ representing the usual single-test significance level, hereby set to the value $t=0.05$: since $H_{i}$ is the hypothesis that the distribution of dyadic motifs established by the $i$-th pair of nodes follows a binomial/Poisson-binomial, rejecting it amounts to connecting the corresponding nodes in the projection. Notice that |H|=N(N1)/2|H|=N(N-1)/2.
@@ VI Benchmarks for the analysis of multivariate time series
To evaluate the statistical significance of the similarity between any two series $i$ and $j$, a benchmark is required. A natural choice is to adopt the class of null models known as Exponential Random Graphs (ERGs). Within such a framework, the constrained maximisation of Shannon entropy
@@
leads to assign the generic biadjacency matrix $\mathbf{B}$ the probability
@@
whose value is determined by the vector $\bm{C}(\mathbf{B})$ of topological constraints via the Hamiltonian $H(\bm{\theta},\bm{C}(\mathbf{B}))=\bm{\theta}\cdot\bm{C}(\mathbf{B})=\sum_{i}\theta_{i}C_{i}(\mathbf{B})$. The use of linear constraints allows $P(\mathbf{B})$ to be written in a factorised form, i.e. as the product of pair-specific probability coefficients gallo2025statistically.
@@
Let us now generalise such a framework to accommodate models for studying binary multivariate time series. To this aim, we will follow Livan2020; gallo2024testing, which, in turn, builds upon the analytical approach introduced in Park2004 and further developed in Squartini2011.
Let us, now, make a general observation: as each of the $N$ time series is defined by $T$ time steps, in corresponding of which either the value 1-1 or the value +1+1 can be assumed, the ensemble is constituted by all possible $|\mathbb{B}|=2^{N\cdot T}$ binary, signed biadjacency matrices.
@@ VI.1 Signed Random Graph Model
i.e. by the two global constraints $B^{+}(\mathbf{B})$ and $B^{-}(\mathbf{B})$. Although being formally analogous to the BiSRGM gallo2024testing, we will refer to it as bSRGM.
According to the bSRGM, each element of a multivariate time series is a random variable whose behaviour is described by the coefficients
@@
In order to employ the bSRGM for studying real-world multivariate time series, the parameters that define it need to be properly tuned: more specifically, one needs to ensure
@@
with the symbol $\mathbf{B}^{*}$ indicating the empirical multivariate time series under analysis. To this aim, one can maximise the likelihood function $\mathcal{L}_{\text{bSRGM}}(\alpha,\gamma)\equiv\ln P_{\text{bSRGM}}(\mathbf{B}^{*}|\alpha,\gamma)$, where $P_{\text{bSRGM}}(\mathbf{B}^{*}|\alpha,\gamma)$ is defined as in eq. 112, with respect to the unknown parameters that define it Garlaschelli2008. Such a recipe leads us to find
@@ VI.2 Signed Configuration Model
i.e. by the $2(N+T)$ local constraints $\{k_{i}^{+}(\mathbf{B})\}_{i=1}^{N}$, $\{k_{i}^{-}(\mathbf{B})\}_{i=1}^{N}$, $\{\kappa_{t}^{+}(\mathbf{B})\}_{t=1}^{T}$ and $\{\kappa_{t}^{-}(\mathbf{B})\}_{t=1}^{T}$. Although being formally analogous to the BiSCM gallo2024testing, we refer to it as bSCM.
According to the bSCM, each element of a multivariate time series is a random variable whose behaviour is described by the coefficients
@@
in other words, $b_{it}$ obeys a generalised Bernoulli distribution whose probability coefficients are determined by the (Lagrange multipliers of the) imposed constraints: the series $i$, at time $t$ assumes the value +1+1 with probability $p_{it}^{+}$ and the value 1-1 with probability $p_{it}^{-}$.
Figure 3: Consistency checks regarding the distribution of the signature for subject $\bm{\#100307}$. Comparison between the distribution of the dyadic signature induced by the bSRGM/bSCM (left panel/middle panel) and their numerical counterparts, obtained by explicitly sampling $10^{5}$ realisations from the corresponding ensemble: in both cases, the KS test rejects the hypothesis that the two distributions coincide. The box-plots summing up the distributions of the KS-scores for the bSRGM and the bSCM, obtained by explicitly sampling $10^{3}$ realisations from the corresponding ensemble (right panel), show that, under both models, the median KS-score amounts to $\text{KS}_{5\%}\simeq 0.965$, the $95\%$ of the values ranging between the $2.5$-th and the $97.5$-th percentiles reading $q_{2.5},q_{97.5}]=[0.961,0.970]$. In words, our sampling procedure can be considered satisfactory enough to reproduce the analytical distributions defined by eq. 44 and eq. 50, even in case of statistical disagreement.
@@
let us maximise the likelihood function $\mathcal{L}_{\text{bSCM}}(\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})\equiv\ln P_{\text{bSCM}}(\mathbf{B}^{*}|\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})$, where $P_{\text{bSCM}}(\mathbf{B}^{*}|\bm{\alpha},\bm{\gamma},\bm{\delta},\bm{\eta})$ is defined as in eq. A, with respect to the unknown parameters that define it Garlaschelli2008 (see also Appendix A). Such a recipe leads us to find
@@
a system that can be solved only numerically, along the guidelines provided in vallarano2021 (see also Appendix B).
Figure 4: Distributions of the signed, bipartite, and monopartite connectance. Left panel: distributions of the positive and negative bipartite connectance, respectively defined as $c^{+}=B^{+}/(N\cdot T)$ and $c^{-}=B^{-}/(N\cdot T)$, across our $100$ subjects. Middle and right panel: distributions of the positive and negative monopartite connectance, respectively defined as $\rho^{+}=2L^{+}/N(N-1)$ and $\rho^{-}=2L^{-}/N(N-1)$, across the three different projections of our $100$ subjects. Bin widths are computed according to the Freedman-Diaconis rule. Averages are plotted as vertical, dashed lines. How can a large number of bipartite, negative links co-exist with the large number of positive links characterising the naïve projection? Bipartite, negative links are evidently arranged into concordant motifs that, in turn, let the naïve and bSRGM-induced projections be populated by a majority of positive links. Employing the bSCM mitigates this situation, as the number of positive and negative links is now rebalanced.
@@ VII Dataset description
We have applied the methodology above to 3T rs-fMRI data from $100$ unrelated subjects of the Young Adult Human Connectome Project (HCP) VanEssen2013.
Overall, this dataset includes $1200$ participants aged $22$ to $35$ of whom $1087$ underwent at least one rs-fMRI scan. Spontaneous, slowly fluctuating brain activity was measured during fMRI, using a Multi-Band Accelerated Echo-Planar-Imaging (EPI) 2D BOLD sequence with whole-brain coverage, while subjects maintained fixation on a central cross and were instructed to lie still and rest quietly - $\text{TR}=720$ ms, $\text{TE}=33.1$ ms, $\text{flip angle}=52$ deg, $\text{FOV}=208\times 180$ $m$ ($\text{RO}\times\text{PE}$), $\text{matrix}=104\times 90$ ($\text{RO}\times\text{PE}$), $\text{slice thickness}=2.0$ $m$, $\text{number of slices}=72$, $\text{resolution}=2.0$ mm isotropic voxels, $\text{multiband factor}=8$, $\text{echo spacing}=0.58$ ms, $\text{BW}=2290$ Hz/Px, $\text{frames per run}=1200$, $\text{run duration\>(min:sec)}=14:33$. Within each session, oblique axial acquisitions alternated between phase encoding in a right-to-left (RL) direction in one run and phase encoding in a left-to-right (LR) direction in the other run.
fMRI recordings underwent a two-step pre-processing procedure. First, the data were motion-corrected, intensity-adjusted, and normalised Glasser2013. Afterwards, structured noise was corrected using the FIX denoising pipeline Griffanti2014. This clean-up included $24$ confound time series derived from the motion estimation (the $6$ rigid-body parameter time series, their backwards-looking temporal derivatives and all $12$ resulting regressors squared Satterthwaite2013). On the motion parameters, the temporal high-pass filtering was previously applied. They are, then, regressed out of the data aggressively, as they are not expected to contain variance of interest. To improve the length of BOLD time series, the right-to-left and left-to-right runs were concatenated after the mean across time removal. Finally, each brain was parcellated into $116$ regions of interest (ROIs) over which the signal was averaged. More specifically, the cortical regions were parcellated following the Schaefer $100$ atlas Schaefer2018 and the subcortical ones following the Tian $16$ atlas Tian2020. As a final result, we obtained a collection of $100$ sets of time series, each one having dimensions $N\times T=116\times 2400$.
@@ VIII Results
Let us, now, employ our validation procedure to carry out a ‘binary-to-binary’ analysis, i.e. obtain statistically validated binary projections of binary multivariate time series, defined by the ‘rule’ that any two nodes sharing a significantly large number of concordant (discordant) motifs are connected by a positive (negative) edge.